OpenStudy (fairy733189):

A quadratic equation is shown below: x^2 - 14x + 41 = 0 Which of the following is the first correct step to write the above equation in the form (x - p)^2 = q, where p and q are integers? Add 8 to both sides of the equation Add 9 to both sides of the equation Subtract 8 from both sides of the equation Subtract 9 from both sides of the equation

OpenStudy (fairy733189):

@kky10997

OpenStudy (kky10997):

option 1st is correct

OpenStudy (fairy733189):

but how do you figure it out?

OpenStudy (zzr0ck3r):

very good question... as this is not a place for answers...

OpenStudy (kky10997):

its easy

OpenStudy (fairy733189):

then can you explain?

OpenStudy (zzr0ck3r):

its easy if you know how....

OpenStudy (kky10997):

see (x-p)^2=q x^2 +2px+p^2 =q now compare with x^2 -14x +41 with above

OpenStudy (zzr0ck3r):

-2px

OpenStudy (kky10997):

sorry -2px

OpenStudy (fairy733189):

I still dont really understand it

OpenStudy (kky10997):

now we get p= 7 so (x-7)^2= x^2 -14x+49 so to make left hand side perter squre we add 8 both side

OpenStudy (fairy733189):

but where did you get the -2px from?

OpenStudy (kky10997):

(x-p)^2=x^2 -2px + p2